OpenStudy (winsteria):

please help! The product of 9 and a number is 12 less than three times that number. What is the number?

OpenStudy (winsteria):

@TheSmartOne @AriPotta

OpenStudy (anonymous):

`The product of 9 and a number` Well, `product` is the result of what operation?

OpenStudy (winsteria):

Multiplication

OpenStudy (anonymous):

Okay so `The product of 9 and a number` would be \(9\times n\).

OpenStudy (anonymous):

`is` this is just equality \(=\)

OpenStudy (anonymous):

`12 less than three times that number` In this case `less` is an operation. It is subtraction.

OpenStudy (anonymous):

`three times that number` will be \(3\times n\) `12 less than three times that number` will be \(3\times n-12\).

OpenStudy (anonymous):

All together:\[ 9\times n = 3\times n - 12 \]

OpenStudy (anonymous):

Can you solve for \(n\)?

OpenStudy (winsteria):

4...?

OpenStudy (winsteria):

Or -4 maybe

OpenStudy (anonymous):

What was your first step in simplifying?

OpenStudy (anonymous):

By the way \(3\times n\) can be written as \(3n\). We can write it as \[ 9n=3n-12 \]

OpenStudy (anonymous):

If you put \(n=-4\), then you get:\[ 9(-4) = 3(-4)-12\\ -36=-12-12\\ -36=-24 \]This is not correct. We should solve for \(n\) using algebra.

OpenStudy (anonymous):

Do you need a hint?

OpenStudy (winsteria):

Yeah

OpenStudy (anonymous):

Well, we want \(n\) to be on only one side, and so we can use subtraction to get rid of \(3n\) from the right side.

OpenStudy (winsteria):

Okay

OpenStudy (anonymous):

Do you know what the result is when you do that?

OpenStudy (winsteria):

No :/

OpenStudy (anonymous):

\[ \underbrace{9n}_{-3n} = \underbrace{3n-12}_{-3n} \]

OpenStudy (anonymous):

Left side:\[ 9n-3n = (9-3)n = 6n \]

OpenStudy (anonymous):

Right side: \[ 3n-12-3n = (3-3)n-12 = 0n-12=-12 \]The \(3n\) goes away as planned.

OpenStudy (anonymous):

So we have\[ 6n=-12 \]

OpenStudy (winsteria):

-2?

OpenStudy (anonymous):

Yes

OpenStudy (winsteria):

Thank you for explaining it to me yay :) @wio